Assume that the government of Japan wants to control deflation and increase aggregate demand. The government then increased the Reserve Ratio from 10% to 20% hoping that a planned deposit of $5,000,000 will be made.

 

A) Calculate the new total money supply that will be generated from the planned deposits.

B) Explain whether the government policy of increasing the reserve ratio is right.


1
Expert's answer
2021-06-28T15:39:01-0400

Change in reserve ratio "= 20-10 = 10%" %

"= \frac{1}{reserve ratio} × planned deposit"


"= \frac{1}{0.1} × 5000000"


"= 10 \u00d7 5000000"


"= 50000000"

ii) The government policy of increasing the reserve ratio will not be right since it will minimize money stock for its citizens, which will force them to operate under credit cost, which is a challenge to sustainability.